# An Incomplete Loop: Deductive, Inductive, and Abductive Learning in Large Language Models

### v2

@@ 2. RQ2:
      The ability to learn through abduction (proposing hypotheses) is generally not related to learning through induction (few-shot learning).
@@ 2 Three Types of Reasoning in Language Models
Reasoning type Analogue in LMs
Citations
Deductive Instruction following
Wei et al. (2021); Sanh et al. (2021)
Inductive In-context learning
Brown et al. (2020)
Abductive Chain-of-thought (one hypothesis)
Wei et al. (2023); Kojima et al. (2022)
Abductive Externally aided instruction inference (many hypotheses)
Qiu et al. (2023); Wang et al. (2023); Zhu et al. (2023); Yang et al. (2024),
Abductive LM-guided instruction inference (many hypotheses)
This paper
@@
We first give a formalized definition of instruction following, in-context learning, and instruction inference, relating these processes to deductive reasoning, inductive reasoning, and abductive reasoning respectively. Figure 1 gives a schematic of the reasoning loop, while Table 1 gives examples of work falling in each category. [^2] We note that there is some disagreement in the philosophy literature about the exact distinction about inductive and abductive reasoning. Here we use commonly-cited definitions, which happen to distinguish in-context learning from the more explicit process of verbalizing and testing hypotheses. Throughout this section, we assume that we have an input $x$, and we want to produce an output $y$ with an autoregressive LM conditioned on some additional piece of data $D$ that specifies the target task: $p_{LM}(y\mid x,D)$. We use in-context learning of linear functions as a running example.
@@ 3 Methods
P(data) We use a separate LM, 444text-davinci-002 was used as the logprobs-generating LM for all models to generate log probabilities for in-context examples given the hypothesis. $\texttt{score}_{\text{P(data)}}(h_{i},D_{i})$ is the sum of log-probabilities of tokens of $D_{i}$.
@@ 4 Domains and Evaluation
Domain
In-context examples ($D_{i}$)
Query example ($x_{i}$)
Expected answer ($y_{i}$)
Hypothesis example ($h_{i}$)
Functions
(-10, -213), (9, 167), (4, 67), …
15
287
$f(x)=20x-13$
Colours
(lug dax, blue green), (lug zup, blue yellow), (lug bluf, blue blue), …
lug walm dax bluf
blue blue blue green green
lug $\rightarrow$ blue
Kalamang vocab (ek)
(That is their place., Tompat ma me muin), (I’m getting pandandus, I want to make a mat.), …
Sakina is pouching guavas.
Sakina sarimara lawat.
guava $\rightarrow$ sarim
Kalamang grammar (ek)
Order of subject and verb: SV
@@ 4.1 Linear Functions
Following investigations of language models’ in-context learning abilities (Garg et al., 2023; Akyürek et al., 2023), we construct a dataset of 40 linear functions $f(x)=ax+b$, where $a$ and $b$ are uniformly sampled from the integers $[-20,20]$, along with 5 test examples for each function, yielding a 200 test examples. For each test example, we randomly generate 5 in-context examples of the function $(x_{i},y_{i})_{i=1}^{5}$ and one test example. The test example and inputs for in-context examples are also uniformly sampled from the integers $[-20,20]$. We refer to this as the functions domain. Prompts used are given in Appendix B.
@@ 4.2 Simple artificial languages
Inspired by compositional instruction-following datasets, we generate a simple dataset where the inputs are nonce words such as lug, and the outputs are colour terms such as blue (Lake & Baroni, 2017; Lake et al., 2019). We call the expanded dataset the colours domain. The ruleset for this domain can be found in Appendix E. We generate 200 test examples and 800 training examples. [^6] We generate this dataset instead of using the original miniscan to mitigate memorization of the original nonce words, as well as to gain more test examples, as the original dataset has only 10. Prompts used are in Appendix F.
Hypotheses Following the original miniscan, we create a fixed minimal set of 5 in-context examples that contains each nonce word at least once, and from which the meaning of each nonce word can be reasonably inferred. During instruction inference, we isolate one nonce word at a time from the sentence, and have the model try to induce the vocabulary mapping for that word from 5 retrieved in-context examples containing that word. The ground truth grammar was written by the author, and consisted of production rules for each nonce term. Prompts used in the abductive and base tasks can be found in Appendix F
@@ 5 Results
In this section, we first evaluate models’ concrete performance across different domains (RQ1:). We highlight the significant improvement instruction inference offers in some cases in synthetic tasks, yet despite this, improvements are not uniform across tasks, and not attested in challenging domains like Kalamang. Our analysis also highlights an inconsistent correlation between the quality of generated hypotheses and few-shot learning success (RQ2:), meaning that the ability to generate or follow instructions doesn’t reliably predict task mastery or vice versa. These results (Table 3) suggest that while structured instructions can boost performance in simpler scenarios, their impact is less predictable in complex settings. Furthermore, the relationship between instruction induction, instruction following, and in-context learning is complex, and each capability may rely on separate unknown aspects of model architecture, training procedures, or data.
Domain
Deductive reasoning works on average
Abductive reasoning works on average
Hypothesis proposal works on average
Abductive reasoning related to inductive reasoning
@@ 5.1 When does instruction inference improve over in-context learning?
How effective are models’ induced instructions? Self-generated instructions also improve on the baseline in many cases, with variations by domain. In linear functions, models’ hypothesis induction markedly surpasses the few-shot baseline, with both the verbalized confidence and log-probability based reranking methods yielding comparable improvements (see Figure 3(A) again). For the colours domain (3(B)), instruction inference benefits performance, though not as strongly as in linear functions. chrF scores follow a similar trend, and are depicted in Appendix G. Interestingly, for gpt-4-turbo and Llama-2-7b, using the models’ self-proposed hypotheses benefits performance more than using the ground-truth grammar for the colours language, despite the fact that self-proposed are not always correct. Unlike the two synthetic domains, inducing grammar and vocabulary items for Kalamang does not improve translation metrics in most cases.
@@ 5.2 How does the ability to induce instructions relate to in-context learning?
GPT-3.5-turbo 0.2685 $5.013\times 10^{-15}$ 0.5570 $2.747\times 10^{-67}$
GPT-4-turbo 0.8485 $2.153\times 10^{-292}$ 0.9132 0.0
Llama-2-7b -0.0069 0.9012 0.1455 0.0081
Llama-2-70b 0.1393 0.0595 -0.0175 0.8222
@@
In linear functions, we examine GPT-3.5-turbo because of GPT-4-turbo’s consistent accuracy in selecting correct hypotheses. [^9] This means agreement cannot be computed with the point-biserial correlation The analysis reveals a chance-level agreement (0.0013), suggesting GPT-3.5-turbo may be able to predict outputs without identifying the underlying function well, or vice versa. This reveals a dissociation between prediction accuracy and instruction induction.
@@ 6 Related Work
Hypothesis Proposal with Language Models Recent work explores hypothesis proposalto improve language model performance in synthetic tasks, namely ACRE, the original MiniSCAN, ListOps, and versions of the ARC dataset (Qiu et al., 2023; Wang et al., 2023). These methods often rely on domain-specific interpreters or code generation by LMs, akin to our functions domain’s ground-truth reranker. We additionally explores probability-based reranking for rule selection across different domains and assesses the accuracy of model-induced rules prior to reranking.
@@
Program Synthesis and Library Learning The general approach we outline for abductive reasoning can be considered a soft form of program synthesis with language models. Library learning, in which a set of reusable tools is learned from examples, is related (Ellis et al., 2020). A similar approach to ours using LLMs is proposed by Zhu et al. (2023), who use a similar two-step process to learn a library of rules for an arithmetic domain as well as a previous synthetic dataset (Sinha et al., 2019). We find similar gains on synthetic domains, and further examine challenges in applying abductive reasoning in complex domains.
LM-generated Instructions Instruction backtranslation is a related concept (Honovich et al., 2022; Li et al., 2024), in which an LM generates instructions for textual data. However, it differs in that the proposed instructions are not directly used at the same time that they are proposed, and it does not focus on generating rules.
@@ 7 Conclusion
We have examined the interplay between deductive, inductive, and abductive reasoning in LMs through the tasks of hypothesis proposal, in-context learning, and self-generated instruction following. Across three domains (linear function learning, artificial language translation, and Kalamang translation), we show that instruction inference is able to improve over few-shot prompting in simple synthetic domains, but that the relationship between these types of reasoning is complex, and they may not work together as expected presently when models are solving complex tasks. As abductive reasoning seems to be a relatively weaker capability in current language models as compared to instruction following, future work could develop more advanced mechanisms for natural-language hypothesis verification and correction. Enhancements in these areas could accelerate progress towards models capable of autonomous learning and self-improvement without the need for fine-tuning.
@@ Appendix B Prompts for Linear Functions Domain
Prompt Type
Usage
Prompt Text
Base system prompt
For reasoning with in-context examples
@@
You are a problem solving system. Your job is to use the input-output pairs to solve the problem as well as you can.
Hypothesis proposal system prompt
For proposing hypotheses based on in-context examples
@@
You are a pattern recognition system. Your job is to come up with a function that describes the data as well as you can.
Instruction following system prompt
For applying a proposed hypothesis or ground-truth hypothesis to the input
@@
You are a problem solving system. Your job is to apply the function to the data in order to produce an answer.
Few-shot examples prompt
For reasoning with in-context examples only
Return the output preceded by ’Output:’ Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Input: {query input}
Prompt with ground-truth hypothesis
Used when prompting the model to directly apply the correct hypothesis to the input. In-context examples are also included.
@@
Function: {The real function} Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Return the output preceded by ’Output:’ Input: {query input}
Prompt for hypothesis induction
Used to have the model propose a single hypothesis for the function based on in-context examples
@@
Write the function that captures the relationship between inputs and outputs. You should write it in the form y = ax^0 + bx^1. Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Function (please write explicitly in the exact form" ‘Output: y = ax^0 + bx^1)’:
Prompt with a self-induced hypothesis
Used similarly to the ”prompt with ground-truth hypothesis”, except with a self-generated hypothesis. The wording is slightly changed.
@@
Use this function to apply to the input example to get the correct output. { model’s hypothesis } However, just write the output like what’s shown in these examples. Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Return the output preceded by ’Output:’ Input: {query input}
Prompt for zero-shot chain-of-thought
Used to encourage the model to generate a chain of thought.
@@
Return the output preceded by ’Final Output:’ Input: {input1} Output: {output1} Input: {input2} Output: {output2} Let’s think step by step about what the function could be. Remember to write down ’Final Output:’ before your final answer. Input: {query input}
Prompt for hypothesis probability estimate
Used to prompt a language model directly for reranking hypotheses
@@
How likely is this hypothesis about the function to be true given the data? Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Function explanation: {model’s hypothesis} Please give a probability between 0 and 1 inclusive, and only answer with a number. Probability:
Prompt for data logprobs given hyp estimate
Used to rerank hypotheses based on logprobs.
@@
These are examples of applying this function: {model’s hypothesis} Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ...
@@ Appendix F Prompts for Colours Domain
Prompt Type
Usage
Prompt Text
Base system prompt
For reasoning with in-context examples
@@
You are a problem solving system. Your job is to use the input-output pairs to solve the problem as well as you can.’’
Hypothesis proposal system prompt
For proposing hypotheses based on in-context examples
@@
You are a rule induction system. Your job is to figure out the rules underlying a problem and report on them. Use the examples to guide your thinking.
Instruction following system prompt
For applying a proposed hypothesis or ground-truth hypothesis to the input
@@
You are a parser. Carefully use the grammar to parse inputs to determine the correct output.
Few-shot examples prompt
For reasoning with in-context examples only
@@
Return the output preceded by ’Output:’ Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Input: {query input} Remember to start your answer with ’Output:’
Prompt with ground-truth hypothesis
@@
Used when prompting the model to directly apply the correct hypothesis to the input. In-context examples are also included.
Use this grammar to parse the input example to get the correct output. Grammar: {colours domain grammar} Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Return the output preceded by ’Output:’ Input: {query input}
Prompt for hypothesis induction
Used to have the model propose a single hypothesis for the translation of a word.
@@
The below examples contain the nonce word {word}. Using the examples, deduce what {word} means. Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Write your answer like this: word -> meaning. Meaning can be a word or a general rule dependent on the context. Rule:
Prompt with a self-induced hypothesis
Used similarly to the ”prompt with ground-truth hypothesis”, except with a self-generated hypothesis. The wording is slightly changed.
@@
Use this grammar to parse the input example to get the correct output. { model’s hypothesis grammar } However, just write the output like what’s shown in these examples. Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Return the output preceded by ’Output:’ Input: {query input}
Prompt for zero-shot chain-of-thought
Used to encourage the model to generate a chain of thought.
@@
Return the output preceded by ’Final Output:’ Input: {input1} Output: {output1} Input: {input2} Output: {output2} Let’s think step by step about what the translation could be. Work through your answer step by step and show your work. Remember to write down ’Final Output:’ before your final answer. Input: {query input}
Prompt for hypothesis probability estimate
Used to prompt a language model directly for reranking hypotheses
@@
These are examples of the translation of the word {word}. Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Given these examples, how likely is this hypothesis about the meaning of {word}? {model’s word hypothesis} Please give a probability between 0 and 1 inclusive, and only answer with a number. Probability:
Prompt for data logprobs given hyp estimate
Used to rerank hypotheses based on logprobs.
@@
These are examples of applying this function: {model’s word hypothesis} Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ...

### v3

# An Incomplete Loop: Instruction Inference, Instruction Following, and In-context Learning in Language Models

@@ 2. RQ2
      The ability to learn through abduction (proposing hypotheses) is generally not related to learning through induction (few-shot learning). Deductive reasoning is generally strong, with large performance gains over using inductive reasoning alone when the provided hypothesis is correct.
@@ 2 Three Types of Reasoning in Language Models
Reasoning type Analogue in LMs Citations
Deductive Instruction following Wei et al. (2022a); Sanh et al. (2022)
Inductive In-context learning Brown et al. (2020)
Abductive Chain-of-thought (one hypothesis) Wei et al. (2022b); Kojima et al. (2022)
Abductive Externally aided instruction inference (many hypotheses) Qiu et al. (2024); Wang et al. (2024); Zhu et al. (2023); Yang et al. (2024),
Abductive LM-guided instruction inference (many hypotheses) This paper
@@
We first give a formal definition of instruction following, in-context learning, and instruction inference, relating these processes to deductive reasoning, inductive reasoning, and abductive reasoning respectively. (Other ways of interacting with LMs may also evoke one or more of these forms of reasoning, but we focus on important and widely used prompting strategies.) Figure 1 gives a schematic of the reasoning loop, while Table 1 gives examples of work falling in each category. [^2] We note that there is some disagreement in the philosophy literature about the exact distinction about inductive and abductive reasoning. Here we use commonly-cited definitions, which happen to distinguish in-context learning from the more explicit process of verbalizing and testing hypotheses. Throughout this section, we assume that we have an input $x$, and we want to produce an output $y$ with an autoregressive LM conditioned on some additional piece of data $D$ that specifies the target task: $p_{LM}(y\mid x,D)$. We use in-context learning of linear functions as a running example.
@@ 3 Methods
P(data) We use a separate LM, 444text-davinci-002 was used as the logprobs-generating LM for all models. to generate log probabilities for in-context examples given the hypothesis. $\texttt{score}_{\text{P(data)}}(h_{i},D_{i})$ is the sum of log-probabilities of tokens of $D_{i}$.
@@ 4 Domains and Evaluation
Domain In-context examples ($D_{i}$) Query example ($x_{i}$) Expected answer ($y_{i}$) Hypothesis example ($h_{i}$)
Functions (-10, -213), (9, 167), (4, 67), … 15 287 $f(x)=20x-13$
Colours (lug dax, blue green), (lug zup, blue yellow), (lug bluf, blue blue), … lug walm dax bluf blue blue blue green green lug $\rightarrow$ blue
Kalamang vocab (ek) (That is their place., Tompat ma me muin), (I’m getting pandandus, I want to make a mat.), … Sakina is pouching guavas. Sakina sarimara lawat. guava $\rightarrow$ sarim
Kalamang grammar (ek) “” Order of subject and verb: SV
@@ 4.1 Linear Functions
Following investigations of language models’ in-context learning abilities (Garg et al., 2022; Akyürek et al., 2023), we construct a dataset of 40 linear functions $f(x)=ax+b$, where $a$ and $b$ are uniformly sampled from the integers $[-20,20]$, along with 5 test examples for each function, yielding 200 test examples. For each test example, we randomly generate 5 in-context examples of the function $(x_{i},y_{i})_{i=1}^{5}$ and one test example. The test example and inputs for in-context examples are also uniformly sampled from the integers $[-20,20]$. We refer to this as the functions domain. Prompts used are given in Appendix B.
@@ 4.2 Simple Artificial Languages
Inspired by compositional instruction-following datasets, we generate a simple dataset where the inputs are nonce words such as lug, and the outputs are colour terms such as blue (Lake & Baroni, 2018; Lake et al., 2019). We call the expanded dataset the colours domain. The ruleset for this domain can be found in Appendix E. We generate 200 test examples and 800 training examples. [^6] We generate this dataset instead of using the original miniscan to mitigate memorization of the original nonce words, as well as to gain more test examples, as the original dataset has only 10. Prompts used are in Appendix F.
Hypotheses Following the original miniscan, we create a fixed minimal set of 5 in-context examples that contains each nonce word at least once, and from which the meaning of each nonce word can be reasonably inferred. During instruction inference, we isolate one nonce word at a time from the sentence, and have the model try to induce the vocabulary mapping for that word from 5 retrieved in-context examples containing that word. The ground truth grammar was written by the author, and consisted of production rules for each nonce term. Prompts used in the abductive and base tasks can be found in Appendix F.
@@ 5 Results
In this section, we first evaluate models’ concrete performance across different domains (RQ1). We highlight the significant improvement instruction inference offers in some cases in synthetic tasks, yet despite this, improvements are not uniform across tasks, and not attested in challenging domains like Kalamang. Our analysis also highlights an inconsistent correlation between the quality of generated hypotheses and few-shot learning success (RQ2), meaning that the ability to generate or follow instructions doesn’t reliably predict task mastery or vice versa. These results (Table 3) suggest that while structured instructions can boost performance in simpler scenarios, their impact is less predictable in complex settings. Furthermore, the relationship between instruction induction, instruction following, and in-context learning is complex, and each capability may rely on separate unknown aspects of model architecture, training procedures, or data.
Domain Deductive reasoning works on average Abductive reasoning works on average Hypothesis proposal works on average Abductive reasoning related to inductive reasoning
@@ 5.1 When Does Instruction Inference Improve Over In-Context Learning?
How effective are models’ induced instructions? Self-generated instructions also improve on the baseline in many cases, with variations by domain. In linear functions, models’ hypothesis induction markedly surpasses the few-shot baseline, with both the verbalized confidence and log-probability based reranking methods yielding comparable improvements (see Figure 3A again). For the colours domain (3B), instruction inference benefits performance, though not as strongly as in linear functions. chrF scores follow a similar trend, and are depicted in Appendix G. Interestingly, for gpt-4-turbo and Llama-2-7b, using the models’ self-proposed hypotheses benefits performance more than using the ground-truth grammar for the colours language, despite the fact that self-proposed are not always correct. Unlike the two synthetic domains, inducing grammar and vocabulary items for Kalamang does not improve translation metrics in most cases.
@@ 5.2 How Does the Ability to Induce Instructions Relate to In-Context Learning?
GPT-3.5-turbo 0.27 $5.0\times 10^{-15}$ 0.56 $2.7\times 10^{-67}$
GPT-4-turbo 0.85 $2.2\times 10^{-292}$ 0.91 0.0
Llama-2-7b -0.0069 0.90 0.14 0.0081
Llama-2-70b 0.14 0.060 -0.018 0.82
@@
In linear functions, we examine GPT-3.5-turbo because of GPT-4-turbo’s consistent accuracy in selecting correct hypotheses. [^9] This means agreement cannot be computed with the point-biserial correlation. The analysis reveals a chance-level agreement (0.0013), suggesting GPT-3.5-turbo may be able to predict outputs without identifying the underlying function well, or vice versa. This reveals a dissociation between prediction accuracy and instruction induction.
@@ 6 Related Work
Hypothesis Proposal with Language Models Recent work explores hypothesis proposalto improve language model performance in synthetic tasks, namely ACRE, the original MiniSCAN, ListOps, and versions of the ARC dataset (Qiu et al., 2024; Wang et al., 2024). These methods often rely on domain-specific interpreters or code generation by LMs, akin to our functions domain’s ground-truth reranker. We additionally explores probability-based reranking for rule selection across different domains and assesses the accuracy of model-induced rules prior to reranking.
@@
Program Synthesis and Library Learning The general approach we outline for abductive reasoning can be considered a soft form of program synthesis with language models. Library learning, in which a set of reusable tools is learned from examples, is related (Ellis et al., 2023). A similar approach to ours using LLMs is proposed by Zhu et al. (2023), who use a similar two-step process to learn a library of rules for an arithmetic domain as well as a previous synthetic dataset (Sinha et al., 2019). We find similar gains on synthetic domains, and further examine challenges in applying abductive reasoning in complex domains.
LM-generated Instructions Instruction backtranslation is a related concept (Honovich et al., 2023; Li et al., 2024), in which an LM generates instructions for textual data. However, it differs in that the proposed instructions are not directly used at the same time that they are proposed, and it does not focus on generating rules.
@@ 7 Conclusion
We have examined the interplay between deductive, inductive, and abductive reasoning in LMs through the tasks of hypothesis proposal, in-context learning, and self-generated instruction following. Across three domains (linear function learning, artificial language translation, and Kalamang translation), we show that instruction inference is able to improve over few-shot prompting in simple synthetic domains, but that the relationship between these types of reasoning is complex, and they may not work together as expected presently when models are solving complex tasks. As abductive reasoning seems to be a relatively weaker capability in current language models as compared to instruction following, future work could develop more advanced mechanisms for natural-language hypothesis verification and correction. The use of hypothesis proposal during training remains under-explored, and joint training of models on question answering and hypothesis proposal with enforced consistency may help models display more consistent behaviour. Enhancements in these areas could accelerate progress towards models capable of autonomous learning and self-improvement.
@@ Appendix B Prompts for Linear Functions Domain
Prompt Type Usage Prompt Text
@@
Base system prompt For reasoning with in-context examples You are a problem solving system. Your job is to use the input-output pairs to solve the problem as well as you can.
@@
Hypothesis proposal system prompt For proposing hypotheses based on in-context examples You are a pattern recognition system. Your job is to come up with a function that describes the data as well as you can.
@@
Instruction following system prompt For applying a proposed hypothesis or ground-truth hypothesis to the input You are a problem solving system. Your job is to apply the function to the data in order to produce an answer.
@@
Few-shot examples prompt For reasoning with in-context examples only Return the output preceded by ’Output:’ Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Input: {query input}
Prompt with ground-truth hypothesis Used when prompting the model to directly apply the correct hypothesis to the input. In-context examples are also included. Function: {The real function} Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Return the output preceded by ’Output:’ Input: {query input}
@@
Prompt for hypothesis induction Used to have the model propose a single hypothesis for the function based on in-context examples Write the function that captures the relationship between inputs and outputs. You should write it in the form y = ax^0 + bx^1. Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Function (please write explicitly in the exact form" ‘Output: y = ax^0 + bx^1)’:
@@
Prompt with a self-induced hypothesis Used similarly to the ”prompt with ground-truth hypothesis”, except with a self-generated hypothesis. The wording is slightly changed. Use this function to apply to the input example to get the correct output. { model’s hypothesis } However, just write the output like what’s shown in these examples. Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Return the output preceded by ’Output:’ Input: {query input}
@@
Prompt for zero-shot chain-of-thought Used to encourage the model to generate a chain of thought. Return the output preceded by ’Final Output:’ Input: {input1} Output: {output1} Input: {input2} Output: {output2} Let’s think step by step about what the function could be. Remember to write down ’Final Output:’ before your final answer. Input: {query input}
@@
Prompt for hypothesis probability estimate Used to prompt a language model directly for reranking hypotheses How likely is this hypothesis about the function to be true given the data? Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Function explanation: {model’s hypothesis} Please give a probability between 0 and 1 inclusive, and only answer with a number. Probability:
@@
Prompt for data logprobs given hyp estimate Used to rerank hypotheses based on logprobs. These are examples of applying this function: {model’s hypothesis} Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ...
@@ Appendix F Prompts for Colours Domain
Prompt Type Usage Prompt Text
@@
Base system prompt For reasoning with in-context examples You are a problem solving system. Your job is to use the input-output pairs to solve the problem as well as you can.’’
@@
Hypothesis proposal system prompt For proposing hypotheses based on in-context examples You are a rule induction system. Your job is to figure out the rules underlying a problem and report on them. Use the examples to guide your thinking.
@@
Instruction following system prompt For applying a proposed hypothesis or ground-truth hypothesis to the input You are a parser. Carefully use the grammar to parse inputs to determine the correct output.
@@
Few-shot examples prompt For reasoning with in-context examples only Return the output preceded by ’Output:’ Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Input: {query input} Remember to start your answer with ’Output:’
@@
Prompt with ground-truth hypothesis Used when prompting the model to directly apply the correct hypothesis to the input. In-context examples are also included. Use this grammar to parse the input example to get the correct output. Grammar: {colours domain grammar} Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Return the output preceded by ’Output:’ Input: {query input}
@@
Prompt for hypothesis induction Used to have the model propose a single hypothesis for the translation of a word. The below examples contain the nonce word {word}. Using the examples, deduce what {word} means. Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Write your answer like this: word -> meaning. Meaning can be a word or a general rule dependent on the context. Rule:
@@
Prompt with a self-induced hypothesis Used similarly to the ”prompt with ground-truth hypothesis”, except with a self-generated hypothesis. The wording is slightly changed. Use this grammar to parse the input example to get the correct output. { model’s hypothesis grammar } However, just write the output like what’s shown in these examples. Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Return the output preceded by ’Output:’ Input: {query input}
@@
Prompt for zero-shot chain-of-thought Used to encourage the model to generate a chain of thought. Return the output preceded by ’Final Output:’ Input: {input1} Output: {output1} Input: {input2} Output: {output2} Let’s think step by step about what the translation could be. Work through your answer step by step and show your work. Remember to write down ’Final Output:’ before your final answer. Input: {query input}
@@
Prompt for hypothesis probability estimate Used to prompt a language model directly for reranking hypotheses These are examples of the translation of the word {word}. Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ... Given these examples, how likely is this hypothesis about the meaning of {word}? {model’s word hypothesis} Please give a probability between 0 and 1 inclusive, and only answer with a number. Probability:
@@
Prompt for data logprobs given hyp estimate Used to rerank hypotheses based on logprobs. These are examples of applying this function: {model’s word hypothesis} Examples: Input: {input1} Output: {output1} Input: {input2} Output: {output2} ...